OpenStudy (anonymous):

How do I find the exact solutions? (0, 2pi) 3cos x -2= cos 2x

OpenStudy (tkhunny):

I'm thinking \(\cos(2x) = 2\cos^{2}(x) - 1\) might be useful.

OpenStudy (anonymous):

Then would I just set each side to zero?

OpenStudy (tkhunny):

You do not know why you set things to zero, do you? The idea is this, if a*b = 0, then either a = 0 or b = 0 (or both.) If you don't have multiplication, it doesn't do the same thing. If you don't have zero, it doesn't do anything. You need two things multiplied together to make zero. \(3\cos x - 2= \cos 2x\) \(3\cos x - 2= 2\cos^{2}(x) - 1\) \(2\cos^{2}(x) - 3\cos x + 1 = 0\) \((2\cos(x) - 1)(\cos(x) - 1) = 0\) NOW we can start setting things to zero!

OpenStudy (anonymous):

would the solutions end up being pi/3, 5pi/3, pi/3?